Sam and Peter had $42 and $30 respectively. They each spent the same amount of money on a pair of shoes. How much was the pair of shoes if Sam had twice as much money as Peter after buying the pair of shoes?


1
Expert's answer
2022-01-03T16:04:02-0500

Let a pair of shoes cost $x, then after shopping Sam will have 42-x and Peter will have 30-x.

By the condition

(42-x)/(30-x) = 2.

Multiply both sides by (30-x) get

42 - x = 60 - 2x

Moving terms with x to the left and without x to the right, obtained:

x = 18

So the pair of shoes costs $18.